# SpIDER: Spatially Informed Dense Embedding Retrieval for Software Issue Localization

### v1

@@ Abstract
Retrieving code units (e.g., files, classes, functions) that are semantically relevant to a given user query, bug report, or feature request from large codebases is a fundamental challenge for LLM-based coding agents. Agentic approaches typically employ sparse retrieval methods like BM25 or dense embedding strategies to identify relevant units. While embedding-based approaches can outperform BM25 by large margins, they often lack exploration of the codebase and underutilize its underlying graph structure. To address this, we propose SpIDER (Spatially Informed Dense Embedding Retrieval), an enhanced dense retrieval approach that incorporates LLM-based reasoning over auxiliary context obtained through graph-based exploration of the codebase. Empirical results show that SpIDER consistently improves dense retrieval performance across several programming languages.
@@ 1 Introduction
Figure 1: SpIDER workflow. All functions are first ranked by semantic similarity to the issue description, and the top-$K$ functions are retrieved. From these, the top-$C$ functions (where $C\leq K$) serve as centers for spatial exploration along ‘contains’ edges. For each center, we use breath first search to explore neighboring functions within $d$ hops, considering only neighbors that also rank within the top-$N$ by semantic similarity. Then an LLM selects the most relevant neighbors for each center. Finally, selected neighbors are inserted immediately below their corresponding centers in the ranked list.
@@
Recent work has demonstrated the potential of agents powered by Large Language Models (LLMs) to perform automated bug repair and implement basic features within large-scale software repositories (yang2024swe; aider; xia2024agentlessdemystifyingllmbasedsoftware). A critical prerequisite for effective code generation is the precise identification of relevant contextual information, specifically code units such as functions, classes, or files. This code localization task represents a foundational yet notably difficult challenge, particularly when operating at the fine-grained level of function retrieval (jimenez2024swebench; rashid2025swepolybenchmultilanguagebenchmarkrepository; zan2025multiswebenchmultilingualbenchmarkissue; chen2025locagent). Poor code retrieval directly compromises the quality of generated code, leading to incorrect or incomplete fixes, the introduction of new bugs, and substantial increases in both the computational cost and duration of the repair process. Effective code localization is thus essential for realizing the promise of autonomous software engineering agents.
@@
Current approaches to code localization fall into two primary categories. Some recent methods exploit graph representations of code repositories alongside the reasoning capabilities of LLMs (ouyang2025repograph; chen2025locagent; jiang2025cosilsoftwareissuelocalization). These approaches primarily employ sparse retrieval strategies such as BM25 (Robertson1994OkapiATBM25) to identify relevant graph nodes or subgraphs for agent exploration. Other methods train bi-modal encoders that semantically align dense embeddings of code units with issue descriptions(fehr-etal-2025-coret; reddy2025sweranksoftwareissuelocalization). These dense retrieval methods rank code units, such as functions, classes, or files, based on their embedding similarity (typically cosine similarity) to the issue description.
@@
However, existing approaches face three critical limitations. First, conventional contrastive learning-based retrieval methods overlook the structural context of code modules within a repository when ranking them. We observe that in software issue localization, buggy code modules are often spatially proximate within the codebase (e.g., within the same file or in neighboring functions), and top-ranked entities retrieved via dense semantic similarity frequently reside near the actual buggy modules. Consequently, relying exclusively on semantic similarity to select the top-$K$ functions may overlook relevant candidates that are spatially close to highly-ranked modules but exhibit marginally lower embedding similarity scores. Second, existing methods have been primarily evaluated on Python benchmarks (chen2025locagent; fehr-etal-2025-coret), particularly SWEBench-Lite, SWEBench-Verified, and LocBench (jimenez2024swebench; chowdhury2024introducingswebenchverified; chen2025locagent) – leaving their performance on other programming languages largely underexplored. Third, function-level retrieval, which represents the most challenging granularity compared to class-level and file-level retrieval(rashid2025swepolybenchmultilanguagebenchmarkrepository; reddy2025sweranksoftwareissuelocalization; chen2025locagent), remains largely understudied.
@@
To address these limitations, we introduce SpIDER, a novel and simple graph-aware dense retrieval strategy for function-level code localization. SpIDER incorporates spatial locality as a complementary signal to semantic similarity, enabling the joint exploitation of both code content and repository structure within a fixed retrieval budget. Our primary focus is on retrieval methods that can augment downstream tasks, including ranking code modules, precise localization, and automated patch generation for resolving GitHub-style issues. Moreover, to facilitate comprehensive evaluation across multiple programming languages, we also introduce SpIDER-Bench for code localization, encompassing Python, Java, JavaScript, and TypeScript repositories from SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench.
Our primary contributions are as follows:
  • We propose SpIDER, a novel ane simple graph-aware dense retrieval strategy that incorporates both semantic content and code structure to determine the most relevant code functions for a given issue description under a fixed retrieval budget.
@@
  • We introduce SpIDER-Bench, a heterogeneous graph-structured dataset for multi-language code localization. It consists of software issues from existing GitHub repositories in SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench benchmarks, covering Python, Java, JavaScript, and TypeScript, with code graphs that include source code content as node-level features.
  • We provide comprehensive empirical validation of SpIDER against existing dense retrieval methods and the popular sparse retrieval method BM25 across all four languages in SpIDER-Bench.
@@ 2 Related Work
Code Retrieval
Recent approaches like zhang2024codesage; suresh2025cornstack; feng-etal-2020-codebert; guo2021graphcodebertpretrainingcoderepresentations have proposed bi-modal encoder models to encode both code chunks as well as text embeddings in an aligned latent space. However, such models are pretrained on standard natural language to code generation tasks yielding a suboptimal performance for code retrieval based on issue descriptions. fehr-etal-2025-coret; reddy2025sweranksoftwareissuelocalization then finetune such models to the Github-style issues dataset to achieve better embedding alignment between issue descriptions and code representations.
Code Graph Construction
Given a code repository, the goal is to construct a graph that can be utilized by a downstream user (e.g., LLM agent) to understand the code structure in unison with the code content to estimate the likelihood of a module to solve a given issue based on logical reasoning and understanding of the codebase. There are various unique designs of code graphs that can be constructed from a code repository based on the definition and construction of the nodes and edges. For example, in ouyang2025repograph each node of the graph corresponds to a line of code which can of type either definition or reference. The edges between nodes capture the relation between them which can of type ‘invokes’ or ‘contains’. On the other hand, in chen2025locagent, each node corresponds to a code module (function, class, file or directory) which are connected via edges of types ‘contains’, ‘invokes’, ‘imports’ and ‘inherits’. jiang2025cosilsoftwareissuelocalization defines two types of graphs, directory-level graph (called module call graph) and file-level graph (called function call graph). Directory-level graph captures the directory structure where each node is a file which in turn can be expanded into function call graph, each node being a function. The files are connected using ‘imports’ edge whereas functions are connected using ‘invokes’ or ‘inherits’ edges.
@@ 3 Problem Definition and Notations
Given a retrieval budget $K$, issue description $Q$, and a codebase represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $n$ nodes, the task is to retrieve the top-$K$ functions that are most likely to require edits in order to resolve the issue described by $Q$. Here, $\mathcal{V}=\{v_{i}\}_{i=1}^{n}$ is the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. Each node $v_{i}\in\mathcal{V}$ represents a code entity such as a function, class, file, or directory, and entities of interest (in this case function nodes) contain the corresponding source code content. Each edge $e(v_{i},v_{j})\in\mathcal{E}$ denotes a relation between code entities, and can be one of the four edge types, ‘contains‘, ‘imports‘, ‘invokes‘ and ‘inherits‘.
Let $\mathcal{V}^{*}=\{v^{*}_{i}\}^{m}_{i=1}\subseteq\mathcal{V}$ denote the ground-truth set of $m$ relevant function nodes for issue $Q$. Let $\mathcal{V}_{K}=\{\hat{v}_{i}\}^{K}_{i=1}\subseteq\mathcal{V}$ denote the top-$K$ function nodes retrieved by a model. The objective is to maximize Recall@$K$ and Accuracy@$K$ (denoted as Acc@$K$), defined as follows:
$\text{Recall@}K=\frac{|\mathcal{V}^{*}\cap\mathcal{V}_{K}|}{|\mathcal{V}^{*}|}$
$\text{Acc@}K=\mathbb{I}[\mathcal{V}^{*}\subseteq\mathcal{V}_{K}]$
Intuitively, Recall@$K$ measures the proportion of ground truth nodes that appear in the retrieved top-$K$ list. In contrast, Acc@$K$ is a stricter metric: it is 1 for an instance only if all of its ground truth nodes are included in the top-$K$ results. At the dataset level, mean Acc@$K$ reflects the proportion of instances that achieve perfect recall (i.e., Recall@$K$ = 1).
4 Building Code Graphs
To represent a codebase as a graph, we follow chen2025locagent, and define each node as a module that can be either function, class, file or directory, connected by edges of type ‘imports’, ‘invokes’, ‘contains’ and ‘inherits’. We use python AST module [^1] https://docs.python.org/3/library/ast.html for constructing graphs for Python programming language, whereas for Java, JavaScript and TypeScript, we use tree-sitter [^2] https://github.com/tree-sitter implementation to construct similar code graphs. The graph construction starts by navigating the repository. For a directory encounter, it is added as a directory node to the graph. Similarly, if a file is encountered, it is added to the graph if it contains a class or a function, otherwise, it is ignored. Then we analyze the valid files (ending in ‘.py’ for python, ‘.java’ for Java, ‘.js’ for JavaScript and ‘.ts’ for TypeScript) further to handle various types of code lines in the file parsed by the tree-sitter or AST python module.
The core challenge in constructing graphs for Java, JavaScript and TypeScript lies in the many ways functions and classes can be defined and organized. For instance, a class might be declared in one file with no method implementations, while its methods are defined elsewhere, either in a different part of the same file or even in a completely separate file. This separation complicates the process of building a complete and accurate program graph. We explicitly handle diverse scenarios of such declaration/definitions using Tree-sitter. Additional details about the dataset and its construction can be found in section˜A.1.
Table 1: SpIDER-Bench data statistics after graph construction (GT = ground truth)

### v2

@@ Abstract
Retrieving code functions, classes or files that are relevant in order to solve a given user query, bug report or feature request from large codebases is a fundamental challenge for Large Language Model (LLM)-based coding agents. Agentic approaches typically employ sparse retrieval methods like BM25 or dense embedding strategies to identify semantically relevant units. While embedding-based approaches can outperform BM25 by large margins, they often don’t take into consideration the underlying graph-structured characteristics of the codebase. To address this, we propose SpIDER (Spatially Informed Dense Embedding Retrieval), an enhanced dense retrieval approach that integrates LLM-based reasoning along with auxiliary information obtained from graph-based exploration of the codebase. We further introduce SpIDER-Bench, a graph-structured evaluation benchmark curated from SWE-PolyBench, SWEBench-Verified and Multi-SWEBench, spanning codebases from Python, Java, JavaScript and TypeScript programming languages. Empirical results show that SpIDER consistently improves dense retrieval performance by at least $13\%$ across programming languages and benchmarks in SpIDER-Bench .
@@ 1 Introduction
Figure 1: SpIDER workflow. SpIDER augments standard dense embedding retrieval with graph-aware spatial exploration and LLM-based reasoning to improve function-level code localization under a fixed retrieval (K). Given an issue description, all functions are first ranked by semantic similarity, and the top-K candidates are retrieved. The top-C (C<K) functions serve as centers (func 1) for structured neighborhood exploration along contains edges in the code graph, where neighboring functions within d hops are considered if they also rank within the top-N semantically (funcs 2,3,4). An LLM then filters these candidates to identify functions that are likely relevant but under-ranked by semantic similarity alone (func3). The selected neighbors are inserted immediately below their corresponding centers while discarding an equal number of bottom ranked functions in the initial top-K list, thereby increasing coverage of structurally related buggy functions without increasing K. Further algorithmic details are provided in section˜A.7 (algorithm˜1).
@@
Recent work has demonstrated the potential of agents powered by Large Language Models (LLMs) to perform automated bug repair and implement basic features within large-scale software repositories (Yang et al., 2024; Gauthier, 2024; Xia et al., 2024). A critical prerequisite for effective code generation is the precise identification of relevant contextual information, specifically code units such as functions, classes, or files. This code localization task represents a foundational yet notably difficult challenge, particularly when operating at the fine-grained level of function retrieval (Jimenez et al., 2024; Rashid et al., 2025; Zan et al., 2025; Chen et al., 2025). Poor code retrieval directly compromises the quality of generated code, leading to incorrect or incomplete fixes, the introduction of new bugs, and substantial increase in both the computational cost and duration of repair. Thus, effective code localization is essential to realize the promise of autonomous software engineering agents.
@@
Current approaches to code localization fall into two primary categories. Some methods exploit graph representations of code repositories alongside the reasoning capabilities of LLMs (Ouyang et al., 2025; Chen et al., 2025; Jiang et al., 2025). These approaches mainly employ sparse retrieval strategies such as BM25 (Robertson et al., 1994) to identify relevant graph nodes or subgraphs for agent exploration. Other methods train bimodal encoders using contrastive objective to semantically align dense embeddings of code units with issue descriptions (Fehr et al., 2025; Reddy et al., 2025). These dense retrieval methods rank code units based on their embedding similarity to the issue description.
@@
These existing approaches face three critical limitations. First, conventional contrastive learning-based retrieval methods overlook the structural context of code modules within a repository when ranking them (Reddy et al., 2025; Zhang et al., 2024; Suresh et al., 2025). We observe that in software issue localization, buggy code modules are often spatially proximate within the codebase (e.g., within the same file or in neighboring functions), and top-ranked entities retrieved via dense semantic similarity frequently reside near the actual buggy modules. Consequently, relying exclusively on semantic similarity to select the top-$K$ functions may overlook relevant candidates that are spatially close to highly-ranked modules but exhibit marginally lower embedding similarity scores. Second, existing methods have been primarily evaluated on Python (Chen et al., 2025; Fehr et al., 2025), particularly SWEBench-Lite, SWEBench-Verified, and LocBench (Jimenez et al., 2024; Chowdhury et al., 2024b; Chen et al., 2025) – leaving their performance on other programming languages largely underexplored. Third, function-level retrieval, which represents the most challenging granularity compared to class-level and file-level retrieval(Rashid et al., 2025; Reddy et al., 2025; Chen et al., 2025), remains largely understudied.
@@
To address these limitations, we introduce SpIDER, a simple graph-aware dense retrieval strategy for function-level code localization. SpIDER incorporates spatial locality as a complementary signal to semantic similarity, enabling the joint exploitation of both code content and repository structure within a fixed retrieval budget. Our primary focus is on retrieval methods that can augment downstream tasks, including ranking code modules, precise localization, and automated patch generation to resolve GitHub-style issues. To facilitate comprehensive evaluation across multiple programming languages, we also introduce SpIDER-Bench for code localization, encompassing Python, Java, JavaScript, and TypeScript repositories from SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench.
@@
Our primary contributions are as follows: (1) We propose SpIDER, a novel and simple graph-aware dense retrieval strategy that incorporates both semantic content and code structure to determine the most relevant functions for any given issue description under a fixed retrieval budget. (2) We introduce SpIDER-Bench, a heterogeneous graph-structured benchmark for multi-language code localization. It comprises of software issues from existing GitHub repositories in SWEBench-Verified, SWE-PolyBench, and Multi-SWEBench, covering Python, Java, JavaScript, and TypeScript, with code graphs that include source code content as node-level features. We provide comprehensive empirical validation of SpIDER against existing dense retrieval methods and the popular sparse retrieval method BM25 across all four languages in SpIDER-Bench.
@@ 2 Related Work
Code Retrieval and Software Issue Localization  Code retrieval refers to the problem of identifying relevant code locations from a codebase that are responsible for a software issue. Classical approaches are rooted in information retrieval, leveraging lexical or semantic similarity to produce a ranked list of candidate code snippets. Consequently, many existing retrieval systems, including some recent ones, rely on sparse retrievers such as BM25 (Robertson et al., 1994). While BM25 indexing is computationally cheaper than dense embedding generation and vector similarity search, it typically underperforms in retrieval quality when rich semantic representations are available (Reddy et al., 2025; Fehr et al., 2025).
Recent dense retrieval approaches propose bimodal encoder models to encode both code chunks as well as issue embeddings in an aligned latent space (Zhang et al., 2024; Suresh et al., 2025; Feng et al., 2020; Guo et al., 2021). However, these models are typically pretrained on generic natural language–to–code objectives, which leads to suboptimal performance for issue-driven code retrieval. To address this gap, Fehr et al. (2025) and Reddy et al. (2025) introduce bimodal encoders fine-tuned on GitHub-style issue resolution datasets such as SWEBench (Jimenez et al., 2024), resulting in improved alignment between issue and code representations. Despite these gains, SWEBench is limited to Python repositories, and no comparably large multilingual dataset currently exists to support training such encoders across diverse programming languages. Consequently, as real-world codebases span multiple languages, these fine-tuned dense retrievers exhibit degraded performance under domain shifts, reducing their robustness and reliability. We argue that incorporating the commonsense reasoning capabilities of large language models (LLMs), together with structural signals derived from code graphs, introduces more invariant representations that are less sensitive to surface-level language artifacts. Augmenting dense retrievers with these components helps mitigate spurious correlations and improves generalization. We empirically validate this hypothesis on multilingual benchmarks, including SWE-PolyBench (Rashid et al., 2025), Multi-SWEBench (Zan et al., 2025), and SWEBench-Verified (Chowdhury et al., 2024a).
LLM-based retrieval methods  Recent work has shown that large language models (LLMs) are highly effective at localizing and resolving GitHub-style software issues, largely due to their strong reasoning capabilities (Kang et al., 2024; Xia et al., 2024; Luo et al., 2024; Örwall, 2024). In the context of code retrieval, Xia et al. (2024) introduce a simple hierarchical localization strategy driven directly by an LLM, whereas Ouyang et al. (2025) employ an intermediate sparse BM25 retrieval stage to identify relevant files prior to code generation. Similarly, Chen et al. (2025) leverage inverted BM25 indexing to match issue-description keywords with node identifiers or code chunks. Both Ouyang et al. (2025) and Chen et al. (2025) further enable LLM-guided traversal of local subgraphs around the retrieved files or nodes, using edges defined by their respective code graph representations. In contrast, Reddy et al. (2025) focus on re-ranking densely retrieved code chunks using LLMs, with the goal of enhancing semantic retrieval through reasoning. Although this approach improves retrieval performance, it does not exploit the structural signals encoded in code graphs. We aim to bridge this gap by jointly leveraging LLM reasoning, semantic embeddings, and the additional contextual information captured by code graphs, thereby improving overall retrieval effectiveness.
Code Graph Construction  Several recent approaches leverage explicit code graph representations to expose structural information that is otherwise difficult to capture with purely textual or embedding-based methods. The specific utility of such graphs depends critically on how nodes and relations are defined. For example, Ouyang et al. (2025) construct graphs at the line-of-code level, where nodes correspond to variable or module definitions and references, and edges encode invokes and contains relationships. While this design enables fine-grained structural reasoning, it scales poorly with repository size, leading to dense and complex graphs that are challenging for LLM-based agents to traverse at inference time.
To improve scalability, other works adopt coarser graph abstractions. Jiang et al. (2025) propose two complementary graph types: a module call graph, connecting files via imports, and a function call graph, linking functions and classes through invokes and inherits relations. More recently, Chen et al. (2025) introduce a unified graph formulation in which nodes represent functions, classes, files, or directories, and edges capture contains’, invokes’, imports’, and ‘inherits’ relationships. This design strikes a balance between expressivity and tractability, enabling structural reasoning at multiple hierarchical levels while remaining amenable to LLM-guided traversal. Following this line of work, we adopt their graph construction, as it naturally supports retrieval at the file, class, and function levels—granularities that are common across programming languages and well aligned with multilingual code retrieval settings.
@@ 3 Problem Definition and Notations
We formalize the function-level code retrieval task and its evaluation, applicable to any retrieval method operating over structured code representations.
Given a codebase and an issue description $Q$, the goal is to retrieve the top-$K$ functions that are most likely to require edits to resolve the issue. We represent the codebase as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $n$ nodes, where $\mathcal{V}={v_{i}}_{i=1}^{n}$ denotes the set of code entities and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ encodes structural relations. Each node $v\in\mathcal{V}$ corresponds to a code entity, such as a function, class, file, or directory; and edges represent relationships of types contains, invokes, imports, and inherits. Let $\mathcal{V}^{*}={v_{i}^{*}}_{i=1}^{m}\subseteq\mathcal{V}$ denote the ground-truth set of $m$ relevant function nodes for issue $Q$.
Scoring Function  A bi-modal encoder $\mathcal{F}(\cdot)$ maps both code units and issue descriptions to a shared embedding space. The relevance score of node $v$ for issue $Q$ is:
$s_{Q}(v)=\cos\bigl(\mathcal{F}(v),\mathcal{F}(Q)\bigr)$
Baseline Dense Retrieval  Standard dense embedding retrieval (DER) ranks all nodes by their relevance scores and returns the top-$K$ candidates:
$\mathcal{S}_{K}(Q):=\underset{v\in\mathcal{V}}{\mathrm{arg\,top}_{K}}\;s_{Q}(v)$
where $\mathrm{arg\,top}_{K}$ returns $K$ nodes with top-$K$ scores $s_{Q}(\cdot)$. We use notation $\mathcal{S}_{K}$ and $\mathcal{S}_{K}(Q)$ interchangeably.
4 Methodology
Table 1: SpIDER-Bench data statistics after graph construction (GT = ground truth). It summarizes the distribution and granularity of ground-truth edits across languages and datasets, showing that function-level edits dominate across benchmarks while exhibiting substantial variation in edit frequency and localization difficulty across programming languages.